OpenStudy (anonymous):

Find the integral of 1/(2x).

OpenStudy (loser66):

(1/2)lnx +C

OpenStudy (anonymous):

How did you get (1/2)lnx?

OpenStudy (loser66):

\[\int \dfrac{1}{2x}dx=\frac{1}{2}\int \frac{1}{x} dx = \frac{1}{2}lnx +C\]

OpenStudy (loser66):

or you can say it is = \(ln (\sqrt{x})\)

OpenStudy (anonymous):

Thanks.

OpenStudy (loser66):

ok